OpenStudy (anonymous):

A solution has C6H6 and C7H8, and it is 25% C6H6 by mass. The vapour pressures are (at 298K): 94.2 torr for C6H6, and 28.4torr for C7H8. Assuming ideal behaviour, calculate the following: a) P of each sol'n comp's in mixture b) P[tot] above sol'n c) COMPOSITION OF VAPOUR IN MASS PERCENT***

OpenStudy (anonymous):

WAIT! I am going to photocopy my attempt.

OpenStudy (mos1635):

a and b looks OK let me see about c...

OpenStudy (mos1635):

to find percent of each in vapours P1*V=n1*R*T=>...m1=P1*V*Mr1/R*T P2*V=n2*R*T=>...m2=P2*V*Mr2/R*T so (m1/m1+m2)*100%=....... and (m2/m1+m2)*100%

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

But I do not know volume...

OpenStudy (anonymous):

Should I assume that it is constant?

OpenStudy (mos1635):

no it is same for both vapurs and goes away after substitution...

OpenStudy (mos1635):

How did you get that? =equation of ideal gas

OpenStudy (mos1635):

then set n=m/Mr and solve to m